OpenStudy (anonymous):

Simplify. Write each answer in scientific notation. Show your work. (2 x 10^1)(2.21 x 10^3)

myininaya (myininaya):

\[2 \cdot 2.21 =?\] \[10^1 \cdot 10^3=?\]

myininaya (myininaya):

so you know how to multiply 2 and 2.21 right?

OpenStudy (anonymous):

thanks xD

myininaya (myininaya):